OpenStudy (anonymous):

If the perimeter of a rectangle is 56 cm and the length of one side is 12 cm, what is the width of the rectangle?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

16

OpenStudy (anonymous):

sorry i thought i did the math wrong

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

Perimeter \(=2 \times\) (Length +width)

OpenStudy (anonymous):

so 56=2*(12+x), can you solve x from here?

OpenStudy (anonymous):

p=2(l+b); l=12 56=2(12+b) so, b=16

OpenStudy (anonymous):

Thanks everyone.